Re: 485 No Trace - some progress


Robert Simpson
 

Last night tried the Grid Bias adjustment procedure item 10 on page 5-10 and found a problem. The non working scope has no change at the Z Out DC test point when adjusting the A channel intensity. My working 485 changes as expected when adjusting the front panel A channel intensity control. The non working scope stays the same at about 8.2 volts. 8.2 volts is about the same as the working scope with the front panel A channel intensity all the way down. So the beam must be in cutoff condition.

So far have traced to R1703 and the voltage varies OK to that point (same as the working 485). I will continue tracing today.
Bob

Join TekScopes@groups.io to automatically receive all group messages.